# Sparse Multivariate Linear Regression with Strongly Associated Response Variables

### v2

@@ 7.1.3 Results when $(p,q)=(80,80)$
The computing time for ap.MRCE is substantially higher than other procedures, specifically when the tuning parameter associated with the graphical lasso subproblem is close to $10^{-4}$. Given this issue we restricted the tuning parameter set for the $L_{1}$ penalty on the error precision into $\{10^{-2+0.5k}:k=0,1,\ldots,8\}$. We note that this computational setup for ap.MRCE is used for the other simulations where $(p,q)=(80,80)$ in Section 7.2 and H.2 (of the supplementary material).
The complete simulation results are in Figure 2730. The best non-oracle estimator was ap.MRCS. Again ap.MRGCS suffered only when $(s_{1},s_{2})=(0.5,0.1)$, and it showed competitive performance in the other settings. The general performance of ap.MRCE was significantly worse than the previous $(p,q)$ pairs.
@@
In this section, we study the performance of our equicorrelation based estimators under generalized $\Sigma_{*}$ in which $\eta_{i}$’s are now heterogeneous. In H.2 of the supplementary material, we compared our methods to the others in the settings where $\eta_{i}$’s are still heterogeneous but symmetrically distributed. We discovered that MRCS and ap.MRCS showed comparable performance to MRGCS and ap.MRGCS, the two best non-oracle methods in that setting. This may be due to the symmetrically distributed $\eta_{i}$’s ($i\in\{1,\ldots,q\}$) which can smooth out heterogeneous marginal error variances. Now we consider asymmetric cases in this section. We used the same data generating process (16)–(18) except that we considered
@@ 7.2.1 Results when $(p,q)=(20,50)$
The complete results can be found in Figure 3134. MRGCS and ap.MRGCS were the two best non-oracle methods, and consistently outperformed the other methods. Note that MRGCS was slightly better than its approximate counterpart. As opposed to the results shown in H.2.1 of the supplementary material where $\eta_{i}$’s are symmetrically distributed, MRCS and ap.MRCS struggled under the asymmetric setting discussed in this section. With relatively higher $s_{1}s_{2}$ values, ap.MRCE even outperformed these two methods for higher values of $\theta$. TNR/TPR results and trends are similar to those shown in Section 7.1.1.
@@ 7.2.2 Results when $(p,q)=(50,20)$
The complete results are in Figure 3538. ap.MRGCS was the best non-oracle method, although it suffered again when $(s_{1},s_{2})=(0.5,0.1)$. As opposed to the results in $(p,q)=(20,50)$, ap.MRCS was the second-best non-oracle competitor, and it outperformed ap.MRGCS when $s_{2}=0.1$. ap.MRCE showed poor performance generally. Note, the prediction performance of MRCS-Or, ap.MRGCS, and ap.MRCS was substantially better than the others. In TNR, surprisingly, ap.MRGCS was the worst, and ap.MRCS was significantly better than ap.MRGCS. In TPR, ap.MRCE and two lasso methods showed poor performance.
@@ 7.2.3 Results when $(p,q)=(80,80)$
The complete simulation results are in Figure 3942. The results show the best non-oracle estimator in this setting is ap.MRGCS. Again, ap.MRGCS suffered only when $(s_{1},s_{2})=(0.5,0.1)$ where it was outperformed by ap.MRCE as well as sep.lasso (see similar results in H.2.3 of the supplementary material where $\eta_{i}$’s are symmetrically distributed). ap.MRCS was outperformed by ap.MRCE for some values of $\theta$’s, when $(s_{1},s_{2})=(0.1,0.5),\,(0.1,1)$. Note that this can be compared with the results in H.2.3 of the supplementary material where ap.MRCS and ap.MRGCS substantially outperformed ap.MRCE. The relative performance of ap.MRCE was improved. ap.MRCE showed substantially lower TNR compared with the other procedures as $\theta$ gets increased. But, it was the best non-oracle method in TPR.
@@ 8.1 Misspecification of $B_{*}$: When the true $B_{*}$ is non-sparse
In this section, we study how non-sparse true regression coefficient matrix affects the proposed methods. As in previous simulations we measure the prediction and model error for evaluation. We follow the same data generating process (16)–(17). We consider $\eta_{i}$’s in (17) over three different settings: constant $\eta_{i}$’s (Setting A; Section 7.1), heterogeneous and symmetrically distributed $\eta_{i}$’s (Setting B; H.2), heterogeneous and asymmetrically distributed $\eta_{i}$’s (Setting C; Section 7.2). We do not consider different sparsity levels for $B_{*}$. Rather than using (18) for generating regression coefficients matrix $B_{*}$, for each element of $B_{*}$, we generate iid draws from a uniform distribution on $(-1/4,1/4)$ when $(p,q)=(20,50),\,(50,20)$, and on $(-1/10,1/10)$ when $(p,q)=(80,80)$. This regression coefficient matrix is non-sparse, and includes many non-zero, small signals. One can expect that sep.lasso and comb.lasso are likely to suffer from this design. We again set $\theta\in\{0,0.5,0.75,0.9,0.95\}$. In addition to the same set of competitors in Section 7, we also considered combined ridge [c.ridge], and separate ridge [s.ridge]. c.ridge and s.ridge are analogously defined as combined lasso and separate lasso.
The complete results are in Figure 4344. Under the setting A, we denote the sub-settings $(p,q)=(20,50)$, $(50,20)$, $(80,80)$ by A1, A2, and A3, respectively. And, we similarly denote the sub-settings for B and C. MRGCS and ap.MRGCS were the best non-oracle method in all settings, and outperformed ap.MRCE as well as lasso and ridge estimators as $\theta$ increases. We note that ap.MRCE was the third-best non-oracle method. In Setting A, the setting with constant $\eta_{i}$’s, MRCS and ap.MRCS were also regarded as the best non-oracles. Although they showed competitive performance with symmetric and heterogeneous $\eta_{i}$’s when $B_{*}$ has sparsity (see H.2 of the supplementary material), both methods struggled in non-sparse cases (settings B and C). The results from this section indicate that if non-sparse $B_{*}$ has sufficiently small (in absolute value) elements, then there is more gain by joint optimization than lose by misspecified $L_{1}$ penalty.
@@ 8.2 Misspecification of $\Sigma_{*}$: Corrupted compound symmetry to general non-sparse covariance
In this section, we conduct simulation studies to study the effect of misspecification of $\Sigma_{*}$ to our procedure. We are interested in how our equicorrelation based procedures react to a different class of error covariance $\Sigma_{*}$, which is assumed to have a corrupted compound symmetry structure that is defined in (25) below. To generate the data, we used the same procedure that is used throughout Section 7 except that we considered different sparsity level of $B_{*}$ (which will be explained below) and used a corrupted compound symmetry $\Sigma_{*}$ which is computed as follows:
@@
where the columns of $V\in\mathbb{R}^{q\times q}$, $v_{1},\ldots,v_{q}$ are generated by the Gram-Schmidt orthogonalization of a $q\times q$ random matrix whose entries are iid standard normal draws; $D\in\mathbb{R}^{q\times q}$ is a diagonal matrix. For the diagonal entries of $D$, we generated iid draws from a two point distribution that is explained below. Through this we vary the condition number of $\Sigma_{*}$. (25) can be interpreted as a convex combination of a general poorly conditioned matrix and a compound symmetry with variance $0.5$ and correlation $0.9$, which is also poorly conditioned. We considered the corruption level $\omega\in\{0.05,0.5,1\}$. $\omega=1$ leads to a general non-sparse $\Sigma_{*}$ that is completely different than compound symmetry. As $\omega$ decreases, the similarity extent of $\Sigma_{*}$ to the compound symmetry class increases. Denoting Ber$(p,a,b)$ by the generic two point distribution that returns a numeric value $a$ (resp. $b$) with the probability $p$ (resp. $1-p$), the following is the condition numbers according to each setting:
@@
    • $D$ constructed upon iid Ber$(0.5,0.1,10)$ draws: 100 / 189.9075 / 412.7517 (setting A / B / C according to $w=1/0.5/0.05$)
    • $D$ constructed upon iid Ber$(0.4,0.01,10)$ draws: 1000 / 507.617 / 453.8222 (setting D / E / F according to $w=1/0.5/0.05$)
@@
    • $D$ constructed upon iid Ber$(0.2,0.2,10)$ draws: 50 / 64.29671 / 153.7895 (setting G / H / I according to $w=1/0.5/0.05$)
    • $D$ constructed upon iid Ber$(0.6,0.02,10)$ draws: 500 / 265.719 / 186.6711 (setting J / K / L according to $w=1/0.5/0.05$)
Note that a $q\times q$ compound symmetry matrix with $q=50$ (resp. $q=20$) and $\theta=0.9$ has condition number 451 (resp. 181). We considered this setting for the elements of $D$ to avoid the diagonally dominant $\Sigma_{*}$, which can be similar to a diagonal matrix. We present numerical experiments that use the diagonally dominant $\Sigma_{*}$ (where the diagonals of $D$ are generated by a chi-square distribution) in H.3 of the supplementary material. For the sparsity levels of $B_{*}$, we considered $(s_{1},s_{2})\in\{(0.5,0.5),\,(0.5,1),\,(1,1)\}$ where the last pair accounts for dense $B_{*}$. For setting A, we denote the sub-setting of sparsity $(0.5,0.5),\,(0.5,1),\,(1,1)$ by A-1, A-2, and A-3, respectively. We use similar notation to denote the sub-settings from B to L. We consider the same set of competitors methods to those described in Section 8.1.
[…]
@@ 9 Data analysis
We examined a data set from Korea water resources management information system (WAMIS) on the river flow chart of the Han river and its branches. The entire raw data set is public and can be downloaded from the portal (http://www.wamis.go.kr/wkw/flw_dubobsif.do). We used the daily stream flow measured in $m^{3}/s$ throughout the calendar year 2023 from February 1st to December 25th. The 6 selected measurement points (bridges) for the analysis are: Haengju (HJ), Hankang (HK), Gwangjin (GJ), Paldang (PD), Yeojoo (YJ), Nahmhankang (NHK). The first four points are in the main stream of the Han river, and the next two points are in the relatively northern region. Within these 6 points, when considering the average flux in February to April as a reference, flows in YJ, NHK are generally on the same scale (70–100 $m^{3}/s$), but the other 4 points are generally on much higher scale (120–300 $m^{3}/s)$. We set the predictors as the stream flux observations from 1 to 7 days before and the responses as the daily flux at each measurement point. Hence, there are 42 predictors and the leading intercept term, as well as 6 response variables in the model. We used a training set that has 52 observations from February 8th to March 31th. This indicates $(n,p,q)=(52,43,6)$. The test set has 139 observations from August 8th to December 25th. The competitors are MRCE, ap.MRCE, comb.lasso, sep.lasso, MRCS, MRGCS, ap.MRCS, ap.MRGCS. The labels are defined in Section 7. The prediction performance results are summarized in Table 12. For the grand average prediction taken by all the responses (see Table 1), ap.MRCS was the best model, and MRCS was the second-best. MRCE, ap.MRCE, combined lasso followed the next. ap.MRGCS, MRGCS were slightly outperformed by combined lasso but outperformed separate lasso as well as OLS. ap.MRCS also performed the best in three responses (see Table 2), and MRCS was the best in two responses.
@@
Competitors OLS comb.Lasso sep.Lasso MRCS MRGCS ap.MRCS ap.MRGCS MRCE ap.MRCE
River flow 434634.6 265887.3 283362.8 246039.1 268250.1 241223.4 268187.2 261599.1 263000.8
@@
10 Acknowledgement
The authors would like to express gratitude to Professor Aaron Molstad for helpful comments on the proof details. Computational resources were provided by the WVU Research Computing Thorny Flat HPC cluster, which is funded in part by NSF OAC-1726534. Price was partially supported by the National Institute of General Medical Sciences, 5U54GM104942-04. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
@@ Appendix A Derivation of the optimization problems (4)
${\rm det}(\Sigma_{*})$ $={\rm det}\left(\eta^{2}_{*}(1-\theta_{*})\left[I_{q}+\frac{\theta_{*}}{1-\theta_{*}}1_{q}1_{q}^{T}\right]\right)$
$=\eta_{*}^{2q}(1-\theta_{*})^{q}\left(1+q\frac{\theta_{*}}{1-\theta_{*}}\right).$
@@
$\log{\rm det}(\Sigma_{*})$ $=q\log(\eta_{*}^{2})+q\log(1-\theta_{*})+\log\left(1+q\frac{\theta_{*}}{1-\theta_{*}}\right)$
$=q\log(\eta_{*}^{2})+q\log(1-\theta_{*})+\log\left(\frac{1-\theta_{*}+q\theta_{*}}{1-\theta_{*}}\right)$
$=q\log(\eta_{*}^{2})+(q-1)\log(1-\theta_{*})+\log(1+(q-1)\theta_{*}).$
@@
$=(\eta_{*}^{2})^{-1}\left[(1-\theta_{*})^{-1}I_{q}-(1-\theta_{*})^{-1}I_{q}1_{q}(\theta_{*}^{-1}+(1-\theta_{*})^{-1}q)^{-1}1_{q}^{T}(1-\theta)^{-1}I_{q}\right]$
$=(\eta_{*}^{2})^{-1}\left[(1-\theta_{*})^{-1}I_{q}-(1-\theta_{*})^{-2}(\theta^{-1}_{*}+(1-\theta_{*})^{-1}q)^{-1}1_{q}1_{q}^{T}\right]$

### v3

@@ 7.1.3 Results when $(p,q)=(80,80)$
The computing time for ap.MRCE was substantially higher than that of other procedures, particularly when the tuning parameter associated with the graphical lasso subproblem is close to $10^{-4}$. To address this, we restricted the tuning parameter set for the graphical lasso penalty into $\{10^{-2+0.5k}:k=0,1,\ldots,8\}$. This setup was also used for ap.MRCE in the simulations with $(p,q)=(80,80)$ in Section 7.2 and Section 3.2 of the supplementary material [5]. Complete results are provided in Figure 34–37 of the supplementary material. Among the non-oracle estimators, ap.MRCS performed best, while ap.MRGCS only underperformed when $(s_{1},s_{2})=(0.5,0.1)$ and remained competitive otherwise. In contrast, ap.MRCE exhibited significantly poorer performance for this $(p,q)$ setting.
Representative model error comparison plots for the case $(s_{1},s_{2})=(0.5,0.5)$ under the above three $(p,q)$ settings are illustrated in Figure 1(a)1(c).
@@
In this section, we study the performance of our equicorrelation-based estimators under generalized $\Sigma_{*}$ in which $\eta_{i}$’s are now heterogeneous. In Section 3.2 of the supplementary material [5], we compared our methods to the others in the settings where $\eta_{i}$’s are still heterogeneous but symmetrically distributed. We discovered that MRCS and ap.MRCS showed comparable performance to MRGCS and ap.MRGCS, the two best non-oracle methods in that setting. This may be due to the symmetrically distributed $\eta_{i}$’s ($i\in\{1,\ldots,q\}$) which can smooth out heterogeneous marginal error variances. Now we consider asymmetric cases in this section. We used the same data generating process (16)–(18) except that we considered
@@ 7.2.1 Results when $(p,q)=(20,50)$
Complete results are shown in Figures 38–41 of the supplementary material [5]. Among the non-oracle methods, MRGCS and ap.MRGCS performed best, with MRGCS slightly outperforming its approximate version. Unlike the symmetric setting detailed in Section 3.2.1 of the supplementary material [5], MRCS and ap.MRCS performed poorly under the asymmetric setting considered here. For larger $s_{1}s_{2}$ values, ap.MRCE even outperformed these two methods at higher $\theta$ values. The TNR/TPR patterns are similar to those in Section 7.1.1.
@@ 7.2.2 Results when $(p,q)=(50,20)$
Complete results are in Figure 42–45 of the supplementary material [5]. ap.MRGCS was the best non-oracle method, although it suffered again when $(s_{1},s_{2})=(0.5,0.1)$. As opposed to the results in $(p,q)=(20,50)$, ap.MRCS was the second-best non-oracle competitor, and it outperformed ap.MRGCS when $s_{2}=0.1$. ap.MRCE showed poor performance generally. Note, the prediction performance of MRCS-Or, ap.MRGCS, and ap.MRCS was substantially better than the others.
@@ 7.2.3 Results when $(p,q)=(80,80)$
Complete simulation results are in Figures 46–49 of the supplementary material [5]. In this setting, ap.MRGCS was the best non-oracle estimator, except when $(s_{1},s_{2})=(0.5,0.1)$, where it was outperformed by ap.MRCE and sep.lasso (similar to the symmetric case in Section 3.2.3 of the supplementary material [5]). ap.MRCS was also outperformed by ap.MRCE for some $\theta$ values when $(s_{1},s_{2})=(0.1,0.5)$ or $(0.1,1)$.
Representative model error comparison plots for the case $(s_{1},s_{2})=(0.5,0.5)$ under the above three $(p,q)$ settings are illustrated in Figure 1(d)1(f).
@@ 8.1 Misspecification of $B_{*}$: When the true $B_{*}$ is non-sparse
In this section, we study the impact of a non-sparse true regression coefficient matrix on the proposed methods. As in previous simulations, we evaluate performance using prediction and model error under the same data-generating process given in (16)–(17). We consider three settings for $\eta_{i}$’s in (17): constant (Setting A), heterogeneous and symmetric (Setting B), and heterogeneous and asymmetric (Setting C).
Unlike earlier experiments, we do not vary the sparsity of $B_{*}$. Rather than using (18), each element of $B_{*}$ is generated independently from a uniform distribution on (1/4,1/4)(-1/4,1/4) for $(p,q)=(20,50)$ and $(50,20)$, and on (1/10,1/10)(-1/10,1/10) for $(p,q)=(80,80)$. This design yields a non-sparse $B_{*}$ with many small nonzero signals, which is expected to challenge sep.lasso and comb.lasso. We again set $\theta\in{0,0.5,0.75,0.9,0.95}$ and include the same competitors as in Section 7, with the addition of combined ridge [c.ridge] and separate ridge [s.ridge], defined analogously to combined and separate lasso.
Complete results are in Figure 50–51 in the supplementary material [5]. Under Setting A, we label the sub-settings $(p,q)=(20,50)$, $(50,20)$, and $(80,80)$ as A1, A2, and A3, respectively, and use similar labels for Settings B and C. Across all settings, MRGCS and ap.MRGCS were the best non-oracle methods, outperforming ap.MRCE as well as lasso and ridge estimators as $\theta$ increased. ap.MRCE ranked as the third-best non-oracle method. In Setting A, with constant $\eta_{i}$’s, MRCS and ap.MRCS also performed competitively as the best non-oracle methods. However, unlike the sparse case with symmetric and heterogeneous $\eta_{i}$’s (see Section 3.2 of the supplementary material [5]), both methods struggled under non-sparse designs (Settings B and C). The results from this section indicate that if non-sparse $B_{*}$ has sufficiently small (in absolute value) elements, then there is more gain by joint optimization than lose by misspecified $L_{1}$ penalty.
@@ 8.2 Misspecification of $\Sigma_{*}$: Corrupted compound symmetry to general non-sparse covariance
In this section, we conduct simulation studies to study the effect of misspecified $\Sigma_{*}$ on our methods. Specifically, we analyze how our equicorrelation-based methods perform when the true error covariance $\Sigma_{*}$ follows a corrupted compound symmetry structure, defined in (25) below. Data are generated using the same procedure as in Section 7, except that we consider different sparsity levels for $B_{*}$ (described below) and replace $\Sigma_{*}$ with the corrupted compound symmetry structure computed as follows:
@@
where the columns of $V\in\mathbb{R}^{q\times q}$, $v_{1},\ldots,v_{q}$ are generated by the Gram-Schmidt orthogonalization of a $q\times q$ random matrix whose entries are iid standard normal draws; $D\in\mathbb{R}^{q\times q}$ is a diagonal matrix. For the diagonal entries of $D$, we generated iid draws from a two point distribution that is explained below. Through this we vary the condition number of $\Sigma_{*}$. (25) can be interpreted as a convex combination of a general poorly conditioned matrix and a compound symmetry with variance $0.5$ and correlation $0.9$, which is also poorly conditioned. We considered the corruption level $\omega\in\{0.05,0.5,1\}$. $\omega=1$ leads to a general non-sparse $\Sigma_{*}$ that is completely different than compound symmetry. As $\omega$ decreases, the similarity extent of $\Sigma_{*}$ to the compound symmetry class increases. Denoting Ber(p,a,b)(p,a,b) by the generic two point distribution that returns a numeric value $a$ (resp. $b$) with the probability $p$ (resp. 1p1-p), the following is the condition numbers according to each setting:
@@
    • $D$ constructed upon iid Ber(0.5,0.1,10)(0.5,0.1,10) draws: 100 / 189.9075 / 412.7517 (setting A / B / C according to $w=1/0.5/0.05$)
    • $D$ constructed upon iid Ber(0.4,0.01,10)(0.4,0.01,10) draws: 1000 / 507.617 / 453.8222 (setting D / E / F according to $w=1/0.5/0.05$)
@@
    • $D$ constructed upon iid Ber(0.2,0.2,10)(0.2,0.2,10) draws: 50 / 64.29671 / 153.7895 (setting G / H / I according to $w=1/0.5/0.05$)
    • $D$ constructed upon iid Ber(0.6,0.02,10)(0.6,0.02,10) draws: 500 / 265.719 / 186.6711 (setting J / K / L according to $w=1/0.5/0.05$)
Note that a $q\times q$ compound symmetry matrix with $q=50$ (resp. $q=20$) and $\theta=0.9$ has condition number 451 (resp. 181). This setting for $D$ avoids a diagonally dominant $\Sigma_{*}$, that behaves like a diagonal matrix. Additional experiments using diagonally dominant $\Sigma_{*}$ with diagonals of $D$ from a chi-square distribution) are reported in Section 3.3 of the supplementary material [5]. For the sparsity of $B_{*}$, we used $(s_{1},s_{2})\in\{(0.5,0.5),\,(0.5,1),\,(1,1)\}$ where the last pair accounts for dense $B_{*}$. Sub-settings A-1, A-2, and A-3 correspond to these sparsity levels, with analogous notation for settings B–L. The same competitor methods as in Section 8.1 are considered.
Complete results are in Table 5–8 in the supplementary material [5]. In general, the oracle estimator performed best, except in cases A-3, D-3, G-3, H-3, I-3, J-3, K-3, L-3, where sparsity is violated and ridge estimators performed the best. When $B_{*}$ was sparse and $\omega=1$ or $0.5$, our estimators not only performed comparably to lasso methods but also outperformed other competitors in many scenarios. For low corruption levels ($\omega=0.05$), our equicorrelation-based estimators outperformed the others substantially. Even when the true error covariance had varying marginal variances, MRCS and ap.MRCS performed nearly as well as MRGCS and ap.MRGCS and often outperformed the latter two. The results from this section (and Section 3.3 of the supplementary material [5]) indicate that our methods are robust to model misspecification except when $B_{*}$ is nearly dense with large signals and the true error covariance deviates substantially from an equicorrelation.
@@ 9 Data analysis
We examined a data set from Korea water resources management information system (WAMIS) on the river flow chart of the Han river and its branches. The entire raw data set is public and can be downloaded from the portal (http://www.wamis.go.kr/wkw/flw_dubobsif.do). We used the daily stream flow measured in $m^{3}/s$ throughout the calendar year 2023 from February 1st to December 25th. The 6 selected measurement points (bridges) for the analysis are: Haengju (HJ), Hankang (HK), Gwangjin (GJ), Paldang (PD), Yeojoo (YJ), Nahmhankang (NHK). Within these 6 points, when considering the average flux in February to April as a reference, flows in YJ, NHK are generally on the same scale (70–100 $m^{3}/s$), but the other 4 points are generally on much higher scale (120–300 $m^{3}/s)$. We set the predictors as the stream flux observations from 1 to 7 days before and the responses as the daily flux at each measurement point. Hence, there are 42 predictors and the leading intercept term, as well as 6 response variables in the model. We used a training set that has 52 observations from February 8th to March 31th. This indicates $(n,p,q)=(52,43,6)$. The test set has 139 observations from August 8th to December 25th. The competitors are MRCE, ap.MRCE, comb.lasso, sep.lasso, MRCS, MRGCS, ap.MRCS, ap.MRGCS, and the factor model [DrFARM; Zhou et al. [14], Chan et al. [2]] with the number of latent factors equal to one. The labels are defined in Section 7.
Prediction results are summarized in Table 12. For the overall average (Table 1), ap.MRCS performed best, followed by MRCS, with DrFARM, MRCE, ap.MRCE, and combined lasso next. ap.MRGCS and MRGCS slightly underperformed compared to combined lasso but outperformed separate lasso and OLS. By individual responses (Table 2), MRCS was best in two cases, while ap.MRCS led in three.
@@
Competitors OLS comb.Lasso sep.Lasso MRCS MRGCS ap.MRCS ap.MRGCS MRCE ap.MRCE DrFARM
River flow 434634.6 265887.3 283362.8 246039.1 268250.1 241223.4 268187.2 261599.1 263000.8 252807.3
@@
10 Conclusion
We propose a new set of methods for multivariate linear regressions that estimate a dense error covariance matrix in high dimensional settings. As our methods leverage an equi-correlation structure, if modeling assumptions are met, then we expect to see performance gains as the number of responses increases. This is due to the larger number of responses to estimate the common correlation parameter.
Furthermore, in general, our methods are designed to work in both high-dimensional and low-dimensional settings; we recommend using MRCS and MRGS in low-dimensional settings, and we recommend using ap.MRCS and ap.MRGS in high-dimensional settings as they reduce numerical instability, improve computational efficiency, and give accurate estimation of the regression coefficient matrix.
11 Acknowledgment
The authors would like to express gratitude to Professor Aaron Molstad for helpful comments on the proof details. Computational resources were provided by the WVU Research Computing Thorny Flat HPC cluster, which is funded in part by NSF OAC-1726534. Price was partially supported by the National Institute of General Medical Sciences, 5U54GM104942-04, and National Institute of Minority Health and Health Disparities, 5R21MD020187-02. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
@@ Appendix A Derivation of the optimization problems (4)
${\rm det}(\Sigma_{*})$ $=\eta_{*}^{2q}(1-\theta_{*})^{q}\left(1+q\frac{\theta_{*}}{1-\theta_{*}}\right).$
@@
$\log{\rm det}(\Sigma_{*})$ $=q\log(\eta_{*}^{2})+(q-1)\log(1-\theta_{*})+\log(1+(q-1)\theta_{*}).$